++++ INAF LBC+LUCIFER queue observing, 2011 March 29 - 30

Observer: R. Speziali, L. Pentericci, K. Boutsia
Telescope Operator: D. Gonzales
Support Astronomer: O. Kuhn
Instrument: LUCIFER

Summary

Sky partly cloudy at the beginning of the night. Seeing very variable throughout the night
We decide to start with Lucifer but the MOS mechanism failed at the first acquisition
We switch to LBC. When Lucifer gets fixed we switch back to Lucifer
1 hour lost due to clouds. Lucifer has been operational for ~2hr.
At the beginning of the night we lost 4 hours on both instruments because we were warned against using LBC while Lucifer was in an unknown state
When we closed up, Lucifer was not operational

Data Summary

Project Objects Filter ExpTime Notes
Grazian EGS Uspec, r-sloan 2hr seeing ~1.2 for the 1st hour (consindered within specs) and 1.3-1.6'' for the 2nd
Maiolino HDF_K2 200H+K+OrderSep 1hr seeing on the dimm >1.5'' - ARC+FLAT - no Tell because of MOS failure

Details

Begin night with Lucifer

02:09 - Slew to Gallerani SDSSJ1048+4637_ref for acq

02:11 - Error setting MOS position - the mechanism that changes masks is blocked
  • trying to resolve error
  • in the meantime we were advised against moving the telescope or switching to LBC because we could damage Lucifer

02:55 -Ventilation doors closed

04:45 - Dome closed because of clouds

05:48 - Opening dome and swithcing to LBC

06:10 - Pointing correction (WT20_376) - dofpia
  • dimm seeing 1.5''

06:20 - Slew to Grazian EGS_UR_1 - dofpia
  • Red channel - OB execution error / FILTERS currently stopped
  • we restart the OB and now it seems to be working
  • collimated after 3 iterations

06:28 - grazian_EGS_UR_1
  • there is an offset that has no star in the guiding technical chip
  • during this offset the seeing on the blue channel gets worse
  • seeing on Red channel: ~1.00'', seeing on Blue channel ~1.3''

07:09 - dofpia

07:14 - grazian_EGS_UR_2

07:54 - dofpia

07:59 - grazian_EGS_UR_3
  • seeing seems to be rising
  • after allseeing script, this OB seems to be out of specs with average seeing>1.4''

08:39 - dofpia
  • dimm seeing 1.5''-2.0''
  • at this point we could also switch to Lucifer because the MOS problems seems to have been solved
  • we choose to stay with LBC because seeing is out of specs for the Luci programs

08:45 - grazian_EGS_UR_4
  • the image with timestamp lbcb.*.090258* shows a big jump (almost 15'')- all objects doubled
  • also in the lbcb.*.091839* there is a jump that shows as a little tail...
  • when the OB finishes, the T.O is rotating the building to measure the wind that is going up.
  • also this OB seems out of spec - at least the blue side

09:26 - Before doing dofpia we rotate building to measure the wind that seems to go up.
  • since the seeing is getting worse and the grazian data is out of spec we decide to switch to Lucifer

09:37 SWITCH TO LUCIFER

10:08 - Slew to RM7_HDF_K2 for acquisition
  • dark images taken during afternoon 0001-0004
  • dimm seeing 1.6
  • we increased the exposure time (150x10) in order to see the reference sources which are too faint 0005-0009

10:38 - Science acquisition for RM7_HDF_K2 0010-0021
  • we can see the spectrum in 3 slits
  • during the OB the seeing on the dimm is on average ~1.6 with picks of 4.0''

11:48 - Before going to the telluric the germans have to restart the instrument

11:50 - Acquisition script for the telluric 0022-0023

11:59 - MOS mechanism got stuck again while we were acquiring the telluric
  • this time the mask remained in the FPU
  • we tried to modify the acquisition script in order to center the telluric without taking mask to turn out
  • we have to comment the mask and mask ID keywords in the script because the script fails
  • with the blind offset the telluric does not seem to be in any slit 0024-0026
  • we cannot move the masks - we cannot guess the offset - in the meantime we lost guiding because it is too bright

12:40 - Calib Unit In
  • ARC 0027-0030
  • FLAT 0031-0036

12:50 - End of night
  • no DoDarks script because the germans have to work on lucifer
  • at the end of the night we leave lucifer not operational

-- KonstantinaBoutsia - 30 Mar 2011
Topic revision: r9 - 02 Apr 2011, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback